Vaeee:

Find the slope of the line through the points (13,14) and (0,−1). Make sure to simplify fractions and enter your answers with a foward slash (i.e. "/"). If the slope is undefined enter "undefined."

YRJ8498:

Is this a quiz or test?

Vaeee:

practice problem

YRJ8498:

ah okay do you got a screenshot?

Vaeee:

carmelle:

The slope formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] \[(x _{1}, y _{1} ) \rightarrow (13,14)\]\[(x _{2}, y _{2} ) \rightarrow (0,-1)\] \[\frac{ (-1)~-~14 }{ 0-13 }= \frac{ -15 }{ -13 }=\frac{ 15 }{ 13 }\] The slope of the line that passes through those two points is: 15/13

Vaeee:

Thanks babes !!

carmelle:

no problem!! :)